# BaZrS Lights Up: The Interplay of Electrons, Photons, and Phonons in Strongly Luminescent Single Crystals

### v3

@@ Abstract
Chalcogenide perovskites have emerged as a promising class of materials for the next generation of optoelectronic applications, with BaZrS attracting significant attention due to its wide bandgap, earth-abundant composition, and thermal and chemical stability. However, previous studies have consistently reported weak and ambiguous photoluminescence (PL), regardless of synthesis method, raising questions about the intrinsic optoelectronic quality of this compound. In this work, we demonstrate strong, band-to-band-dominated PL at room temperature in high-quality BaZrS single crystals. Despite the narrow, single-component PL emission band, time-resolved PL measurements reveal a carrier lifetime of ns. To understand the origin of the strong PL and short carrier lifetime, we perform multiwavelength excitation and polarization-dependent Raman measurements, supported by first-principles lattice dynamics calculations. We identify all 23 theoretically predicted Raman-active modes and their symmetries, providing a comprehensive reference for future studies. Our results indicate that phonon-assisted carrier decay and strong electron-phonon coupling contribute to the short carrier lifetimes, as evidenced by Raman spectroscopy and DFT calculations. Further studies on compositional variations or partial cation/anion substitutions could mitigate electron-phonon coupling and enhance carrier lifetimes. By establishing a detailed reference for the intrinsic vibrational and optoelectronic properties of BaZrS, this work paves the way for further advancements in chalcogenide perovskites for energy and optoelectronic technologies.
@@ II Methods
We grew single crystals of BaZrS using the chemical vapor transport (CVT) technique. The starting materials – BaS (Sigma Aldrich, 99.9%), Zr (Cerac, 99.8%), and S (Sigma Aldrich, 4N8) – were weighed in stoichiometric proportions for BaZrS and loaded into a glassy carbon boat. The boat was placed inside a quartz ampoule with a diameter of 28 mm and a length of 220 mm. Iodine was added as a transport agent at a concentration of 5 mg/cm. The ampoule was evacuated, sealed, and positioned in a two-zone furnace. The temperature was gradually increased from 950∘C to 1050∘C to initiate the reaction between the starting materials. The ampoule was maintained under a temperature gradient from 1150∘C (hot zone) to 1050∘C (cold zone) for 240 hours, enabling material transport from the hot to the cold end. During this growth process, small cube-shaped crystals formed at the colder end of the ampoule, while an elongated, needle-like crystal, measuring 4 mm in length, grew at the center. After the furnace was cooled to ambient temperature, the ampoule was opened, and volatile gaseous by-products were allowed to evaporate. Finally, the crystal was rinsed with ethanol to remove soluble, moisture-sensitive residues.
The crystal structure of nominally BaZrS was determined using single crystal X-ray diffraction (XRD). Single crystal XRD measurements were performed using a Bruker APEX-II CCD diffractometer with Mo-K radiation ( = 0.71073 Å). Data acquisition was carried out using Bruker APEX3, while cell refinement and data reduction were performed using Bruker SAINT. A numerical absorption correction was applied using Bruker SADABS. Structure solution and refinement was performed with SHELXS97 [40] and SHELXL2018/3.3 [41]. The refinement resulted in a distorted perovskite-type structure with the orthorhombic space group Pnma. The key refinement parameters include the lattice parameters Å, Å, and Å, and , mm, K, , , and 2507 measured reflections (2335 independent reflections with ).
@@
Raman and photoluminescence (PL) spectra were measured using a WITec alpha300 R and Horiba LabRam confocal Raman microscope in a backscattering configuration. Multiwavelength excitation Raman measurements were conducted using 488, 532, 638, and 785 nm lasers. The excitation laser was focused onto the sample surface using a longe-range 50x microscope objective with a numerical aperture of 0.55, yielding beam diameters approximately 1.1 m for the 488 nm laser, 1.2 m for the 532 nm laser, and 1.7 m for the 785 nm laser. To ensure that the selected laser power did not induce heating effects or structural modifications in the sample, a power-dependent study was performed. The Raman spectrum was recorded at a fixed location on the sample while systematically increasing the laser power density, starting from the lowest available setting. Each spectrum was evaluated for shifts in peak positions, broadening of peaks, or the emergence of additional features. The maximum power level at which no such changes were detected was chosen as the optimal power setting for all subsequent measurements. The backscattered light was analyzed using two spectrometers. For 488, 532 and 638 nm excitation, a spectrometer equipped with a holographic grating (150 g/mm for PL, 1800 g/mm for Raman) and a thermoelectrically cooled CCD was used. For measurements with 785 nm excitation, a spectrometer with a 1200 g/mm grating and a deep-depletion cooled CCD was employed. Raman peak positions were calibrated against the Si reference peak at 520 cm. All measurements were conducted at room temperature.
@@ III Results
First, we perform a chemical and structural characterization of the synthesized single crystal. Fig. 1(a) presents an optical microscopy image of the needle-single single crystal of BaZrS. The chemical composition of the crystal was investigated by X-ray fluorescence spectroscopy, showing a Ba/Zr ratio of 0.93, slightly lower than the nominal stoichiometric ratio. The structural analysis of the crystal confirms the distorted-perovskite type, orthorhombic crystal structure with corner-sharing ZrS octahedra (space group Pnma). Fig. 1(b) depicts the determined crystal structure, generated in VESTA from the CIF file obtained through single-crystal XRD structure determination. The obtained lattice parameters are Å, Å, and Å. The atomic position parameters, determined from the structural analysis, are presented in Table 1. These parameters further confirm the precise positions of the constituent elements within the orthorhombic unit cell. The CIF file is available for download in the Supplementary Material.
@@
Figure 2(a) shows the power-dependent PL spectra of the BaZrS single crystal. Previous reports on the PL spectra of BaZrS have been highly inconsistent, often featuring broad, asymmetric, and sometimes even multi-component emission bands [23, 29, 27, 18, 26, 31, 28, 50, 51, 30]. In contrast, we observe a single symmetric emission peak in all our spectra, which can be fitted with a single Gaussian function. The peak width ranges from 0.19 to 0.24 eV at maximum power, which, to the best of our knowledge, is the narrowest emission peak reported for BaZrS. As the width of the PL peak is related to the Urbach energy and the density of shallow defects, this suggests that our BaZrS crystal exhibits higher optoelectronic quality compared to previous reports. It is also worth noting that our measurements were conducted at room temperature, where Ye et al. recently reported no detectable PL from BaZrS single crystals [26]. Finally, we observe no emission features at lower photon energies, where Márquez et al. and Pradhan et al. have reported midgap defect peaks (1.2 eV) attributed to undetected secondary phases and sulfur interstitials, which are hypothesized to quench PL in most BaZrS samples [51, 52].
@@
We observe a shift in the PL emission peak with increasing integration time, as shown in Figure 2(c). While excitonic effects could, in principle, lead to transitions slightly below the bandgap energy, we believe this shift is more likely due to sample heating. Reported PL peak positions in the literature vary widely, ranging from 1.65 eV to 1.95 eV [15]. However, given the strong PL intensity, narrow peak width, and band-to-band-like transition observed in our single crystal, we propose that the peak position of 1.920.01 eV represents the intrinsic bandgap of BaZrS at room temperature. Yuan et al. [57] recently demonstrated that several intrinsic defects form shallow donor levels with relatively low formation enthalpies. In addition to chemical impurities and structural defects, these may contribute to the broad range of reported emission peak positions and peak asymmetries in the literature. The observed peak shift with increasing power density, as well as the influence of total injected energy — illustrated here by the shift with increasing integration time during the data acquisition — underscores the importance of reporting such measurement conditions when comparing PL data across different studies.
@@
Longer TRPL decay times have been observed by Ye et al. with crystals grown on LaAlO [30], and Wei et al. with thin films synthesized by sulfurization of PLD-deposited oxide precursors [23]. However, these longer decay times are believed to originate from carrier localization, likely due to defect trapping, as indicated by biexponential decay behavior. The presence of multiple recombination channels in these samples aligns with the more complex steady-state PL spectra, which often feature several emission peaks [26]. While Surendran et al., Wei et al., and Ye et al. all report biexponential decays, the shorter decay components in their studies are comparable to our characteristic decay of ns. Yuan et al. [57] calculated an upper limit of 33 ns for the non-radiative lifetime in BaZrS synthesized under sulfur-rich conditions. Ye et al. initially reported decay times of up to 50 ns [30], but later clarified that these long lifetimes are likely associated with carrier localization through defect trapping, which explains the coexistence of very long TRPL decay times and low PL quantum yields [26].
@@
Group theory analysis predicts 59 phonon modes for BaZrS (space group Pnma (No. 62), point group mmm) with the following irreducible representations [61]:
@@
A more detailed analysis of the phonon nature in BaZrS can be obtained from the calculated phonon dispersion along high-symmetry directions of the Brillouin zone, as shown in Figure 3(b), along with the elemental phonon density of states (PDOS). Several distinct regions can be identified in the phonon dispersion diagram: (i) the low-frequency region ( cm), predominantly dominated by Ba-related vibrations with smaller contributions from Zr atoms, (ii) the intermediate region (100 – 240 cm), which corresponds to mixed contributions from Zr- and S-related vibrations, and (iii) two high-frequency regions (260 – 280 cm and 370 – 420 cm), primarily attributed to S-related vibrations. Additionally, two phonon gaps are observed: one in the 240 – 260 cm frequency region, and another in the 280 – 370 cm region. These gaps arise from the mass differences between Ba, Zr, and S atoms, as well as the strong bonding within the ZrS octahedra compared to the weaker interactions involving Ba. The phonon gaps restrict available phonon scattering channels, which in turn enhances the energy transfer from excited carriers to the lattice, promoting non-radiative recombination. The existence of phonon gaps in BaZrS suggests enhanced electron-phonon coupling and an increased probability of defect trapping, in agreement with the observations by Ye et al. [26].
@@ V Conclusion
Despite these promising optoelectronic properties, our study also underscores challenges such as short carrier lifetimes and strong electron-phonon interactions, which limit its device performance potential. Future research should explore compositional tuning and defect passivation strategies to mitigate non-radiative recombination losses and enhance the viability of BaZrS for practical applications. By establishing the intrinsic optoelectronic and vibrational properties of BaZrS, this work paves the way for further advancements in chalcogenide perovskites for energy and optoelectronic technologies.
@@ Acknowledgements
The data that support the findings of this study are available from the corresponding authors upon request.

### v4

@@ Abstract
Chalcogenide perovskites have emerged as a promising class of materials for the next generation of optoelectronic applications, with BaZrS attracting significant attention due to its wide bandgap, earth-abundant composition, and thermal and chemical stability. However, previous studies have consistently reported weak and ambiguous photoluminescence (PL), regardless of synthesis method, raising questions about the intrinsic optoelectronic quality of this compound. In this work, we demonstrate strong, band-to-band-dominated PL at room temperature in high-quality BaZrS single crystals, with a PL quantum yield of 0.005%. Despite the narrow, single-component PL emission band, time-resolved PL measurements reveal a carrier lifetime of ns. To understand the origin of the strong PL and short carrier lifetime, we perform multiwavelength excitation and polarization-dependent Raman measurements, supported by first-principles lattice dynamics calculations. We identify all 23 theoretically predicted Raman-active modes and their symmetries, providing a comprehensive reference for future studies. Our results indicate that phonon-assisted carrier decay and nontrivial electron-phonon interactions contribute to the short carrier lifetimes, as evidenced by Raman spectroscopy and DFT calculations. Further studies on compositional variations or partial cation/anion substitutions could mitigate electron-phonon coupling and enhance carrier lifetimes. By establishing a detailed reference for the intrinsic vibrational and optoelectronic properties of BaZrS, this work paves the way for further advancements in chalcogenide perovskites for energy and optoelectronic technologies.
@@ II Methods
We grew single crystals of BaZrS using the chemical vapor transport (CVT) technique. The starting materials – BaS (Sigma Aldrich, 99.9%), Zr (Cerac, 99.8%), and S (Sigma Aldrich, 4N8) – were weighed in stoichiometric proportions for BaZrS and loaded into a glassy carbon boat. The boat was placed inside a quartz ampoule with a diameter of 28 mm and a length of 220 mm. Iodine was added as a transport agent at a concentration of 5 mg/cm. The ampoule was evacuated, sealed, and positioned in a two-zone furnace. The temperature was gradually increased from 950∘C to 1050∘C to initiate the reaction between the starting materials. The ampoule was maintained under a temperature gradient from 1150∘C (hot zone) to 1050∘C (cold zone) for 240 hours, facilitating material transport from the hot to the cold end. During this growth process, an elongated, needle-like crystal, measuring 4 mm in length, grew at the center. After the furnace cooled to ambient temperature, the ampoule was opened, allowing volatile gaseous by-products to evaporate. Finally, the crystal was rinsed with ethanol to remove soluble, moisture-sensitive residues. In contrast to previous studies that employed broader temperature ranges or shorter reaction times – often resulting in polycrystalline material or secondary phases such as BaS, ZrS, or ZrO [18, 15] – we used a more precisely controlled CVT process with a narrower temperature gradient and an extended reaction duration. This optimized protocol enabled the growth of high-quality, phase-pure BaZrS single crystals, as confirmed by XRD and Raman spectroscopy.
The crystal structure of nominally BaZrS was determined using single crystal X-ray diffraction (XRD). Single crystal XRD measurements were performed using a Bruker APEX-II CCD diffractometer with Mo-K radiation ( = 0.71073 Å). Data acquisition was carried out using Bruker APEX3, while cell refinement and data reduction were performed using Bruker SAINT. A multiscan absorption correction was applied using Bruker TWINABS. Structure solution and refinement was performed with SHELXT [40] and SHELXL2018/3.3 [41]. The refinement resulted in a distorted perovskite-type structure with the orthorhombic space group Pnma. The key refinement parameters are Å, Å, Å, Å; ; mm; K; measurement range: , , , ; 16493 observed reflections (for both individua), 2507 reflections after merging in point group mmm (2335 independent reflections with ), completeness: 99.8 %; , ; refined parameters: 30; (all data) = 0.0358, , ; . The crystal was refined as a non-merohedral twin using a HKLF5 file produced by TWINABS with 2 individua of 30 % and 70 % relative intensities.
@@
Raman and photoluminescence (PL) spectra were measured using a WITec alpha300 R and Horiba LabRam confocal Raman microscope in a backscattering configuration. Multiwavelength excitation Raman measurements were conducted using 488, 532, 638, and 785 nm lasers. The excitation laser was focused onto the sample surface using a long-range 50x microscope objective with a numerical aperture of 0.55, yielding beam diameters approximately 1.1 m for the 488 nm laser, 1.2 m for the 532 nm laser, and 1.7 m for the 785 nm laser. To ensure that the selected laser power did not induce heating effects or structural modifications in the sample, a power-dependent study was performed. The Raman spectrum was recorded at a fixed location on the sample while systematically increasing the laser power density, starting from the lowest available setting. Each spectrum was evaluated for shifts in peak positions, broadening of peaks, or the emergence of additional features. The maximum power level at which no such changes were detected was chosen as the optimal power setting for all subsequent measurements. The backscattered light was analyzed using two spectrometers. For 488, 532 and 638 nm excitation, a spectrometer equipped with a holographic grating (150 g/mm for PL, 1800 g/mm for Raman) and a thermoelectrically cooled CCD was used. For measurements with 785 nm excitation, a spectrometer with a 1200 g/mm grating and a deep-depletion cooled CCD was employed. All Raman measurements were performed using a power density of 20 mW/cm, with each spectrum acquired over 60 seconds and averaged over three acquisitions to improve the signal-to-noise ratio. Raman peak positions were calibrated against the Si reference peak at 520 cm. All measurements were conducted at room temperature.
@@ III Results
First, we perform a chemical and structural characterization of the synthesized single crystal. Fig. 1(a) presents an optical microscopy image of the needle-like single crystal of BaZrS. The chemical composition of the crystal was investigated by X-ray fluorescence spectroscopy, revealing a Ba/Zr ratio of . This observation suggests the possibility of slight off-stoichiometry, which may be accommodated by defect formation. However, as detailed later, XRD and Raman spectroscopy confirms the absence of secondary phases, indicating that any deviations from stoichiometry are not associated with macroscopic phase separation. Given the relatively high formation energies of isolated barium vacancies () and cation antisites (, ) predicted by DFT studies [50, 51], these defects are unlikely to exist in significant concentrations, even under barium-poor conditions. Instead, we expect charge-neutral defect complexes – such as or – to form preferentially, as these would maintain charge neutrality and minimize the electrostatic energy of the lattice. Although such complexes have not yet been explicitly investigated in BaZrS through DFT calculations, their formation is strongly supported by the defect chemistry in related chalcogenide materials, such as Cu(In,Ga)S and kesterites [52, 53]. Considering the estimated 4% experimental uncertainty in the XRF measurements, along with our comprehensive structural and optoelectronic characterization presented below, we conclude that the crystal is close to stoichiometric, with any native defects being low in concentration and likely existing as electrically benign complexes.
The structural analysis of the crystal confirms the distorted-perovskite type, orthorhombic crystal structure with corner-sharing ZrS octahedra (space group Pnma). Fig. 1(b) depicts the determined crystal structure, generated in VESTA from the CIF file obtained through single-crystal XRD structure determination. The obtained lattice parameters are Å, Å, and Å. The atomic position parameters, determined from the structural analysis, are presented in Table 1. These parameters further confirm the precise positions of the constituent elements within the orthorhombic unit cell. The CIF has been deposited with the CCDC and assigned CSD Number 2465609.
@@
Figure 2(a) shows the power-dependent PL spectra of the BaZrS single crystal. Previous reports on the PL spectra of BaZrS have been highly inconsistent, often featuring broad, asymmetric, and sometimes even multi-component emission bands [23, 29, 27, 18, 26, 31, 28, 54, 55, 30]. In contrast, we observe a single symmetric emission peak in all our spectra, which can be fitted with a single Gaussian function. The peak width ranges from 0.19 to 0.24 eV at maximum power, which, to the best of our knowledge, is the narrowest emission peak reported for BaZrS. As the width of the PL peak is related to the Urbach energy and the density of shallow defects, this suggests that our BaZrS crystal exhibits higher optoelectronic quality compared to previous reports. It is also worth noting that our measurements were conducted at room temperature, where Ye et al. recently reported no detectable PL from BaZrS single crystals [26]. We observe no emission features at lower photon energies, where Márquez et al. and Pradhan et al. have reported midgap defect peaks (1.2 eV) attributed to undetected secondary phases and sulfur interstitials, which are hypothesized to quench PL in most BaZrS samples [55, 56]. Finally, we estimated the photoluminescence quantum yield (PLQY) to be approximately 0.005% under 532 nm excitation, calibrated against an internal perovskite reference standard with a known PLQY of 0.03% as reported by Lai et al. [57].
@@
We observe a shift in the PL emission peak with increasing integration time, as shown in Figure 2(c). While excitonic effects could, in principle, lead to transitions slightly below the bandgap energy, we believe this shift is more likely due to sample heating. A similar effect is observed when measuring the PL under 488 nm excitation, as shown in Figure S2 in the Supporting Information. Comparison of the PL spectra excited with 532 nm and 488 nm reveals a redshift of approximately 30 meV in the emission peak. While the spectral broadening is negligible, we attribute the redshift to local heating resulting from the smaller penetration depth of the 488 nm laser, as well as to the enhanced contribution of surface states, consistent with observations previously reported in halide perovskite single crystals [62].
Reported PL peak positions in the literature vary widely, ranging from 1.65 eV to 1.95 eV [15]. However, given the strong PL intensity, narrow peak width, and band-to-band-like transition observed in our single crystal, we propose that the peak position of 1.920.01 eV represents the intrinsic bandgap of BaZrS at room temperature. Yuan et al. [50] recently demonstrated that several intrinsic defects form shallow donor levels with relatively low formation enthalpies. In addition to chemical impurities and structural defects, these may contribute to the broad range of reported emission peak positions and peak asymmetries in the literature. The observed peak shift with increasing power density, as well as the influence of total injected energy — illustrated here by the shift with increasing integration time during the data acquisition — underscores the importance of reporting such measurement conditions when comparing PL data across different studies.
@@
Longer TRPL decay times have been observed by Ye et al. using single crystals [30], and Wei et al. with thin films synthesized by sulfurization of PLD-deposited oxide precursors [23]. However, these longer decay times are believed to originate from carrier localization, likely due to defect trapping, as indicated by biexponential decay behavior. The presence of multiple recombination channels in these samples aligns with the more complex steady-state PL spectra, which often feature several emission peaks [26]. While Surendran et al., Wei et al., and Ye et al. all report biexponential decays, the shorter decay components in their studies are comparable to our characteristic decay of ns. Yuan et al. [50] calculated an upper limit of 33 ns for the non-radiative lifetime in BaZrS synthesized under sulfur-rich conditions. Ye et al. initially reported decay times of up to 50 ns [30], but later clarified that these long lifetimes are likely associated with carrier localization through defect trapping, which explains the coexistence of very long TRPL decay times and low PL quantum yields [26].
@@
Group theory analysis predicts 60 phonon modes for BaZrS (space group Pnma (No. 62), point group mmm) with the following irreducible representations [66]:
@@
We note that all observed peaks in the Raman spectra of the single crystal can be attributed to the BaZrS phase, with no evidence for the presence of secondary phases such as ZrS, BaS, BaS, BaZrO, ZrO, or oxysulfides, based on comparison with reference Raman spectra reported in the literature [69, 70, 33].
@@
A more detailed analysis of the phonon nature in BaZrS can be obtained from the calculated phonon dispersion along high-symmetry directions of the Brillouin zone, as shown in Figure 3(b), along with the elemental phonon density of states (PDOS). Several distinct regions can be identified in the phonon dispersion diagram: (i) the low-frequency region ( cm), predominantly dominated by Ba-related vibrations with smaller contributions from Zr atoms, (ii) the intermediate region (100 – 240 cm), which corresponds to mixed contributions from Zr- and S-related vibrations, and (iii) two high-frequency regions (260 – 280 cm and 370 – 420 cm), primarily attributed to S-related vibrations. Additionally, two phonon gaps are observed: one in the 240 – 260 cm frequency region, and another in the 280 – 370 cm region. These gaps arise from the mass differences between Ba, Zr, and S atoms, as well as the strong bonding within the ZrS octahedra compared to the weaker interactions involving Ba. The phonon gaps restrict available phonon scattering channels, which in turn enhances the energy transfer from excited carriers to the lattice, promoting non-radiative recombination. While the existence of phonon gaps in BaZrS suggests enhanced electron-phonon coupling and an increased probability of defect trapping, in agreement with the observations by Ye et al. [26], recent theoretical studies [71] have shown that electron-phonon coupling in BaZrS is notably strong, significantly influencing polaron formation and carrier dynamics. The observed phonon gaps are consistent with a lattice that supports moderate to strong electron-phonon coupling, where specific vibrational modes preferentially interact with charge carriers, leading to polaronic effects. This behavior contrasts with more rigid semiconductors such as GaAs, Si, or GaN, where stronger bonding, lower dielectric polarizability, and higher phonon frequencies suppress such interactions. Therefore, the combination of a soft lattice, a polarizable bonding environment, and the presence of phonon gaps provides evidence for nontrivial electron-phonon interactions in BaZrS, which are crucial for understanding its carrier recombination and transport properties.
@@ V Conclusion
Despite these promising optoelectronic properties, our study also underscores challenges such as short carrier lifetimes and nontrivial electron-phonon interactions, which limit its device performance potential. Future research should explore compositional tuning and defect passivation strategies to mitigate non-radiative recombination losses and enhance the viability of BaZrS for practical applications. By establishing the intrinsic optoelectronic and vibrational properties of BaZrS, this work paves the way for further advancements in chalcogenide perovskites for energy and optoelectronic technologies.
@@ Acknowledgements
The data that support the findings of this study are available from the corresponding authors upon request. The crystallographic information file (CIF) has been deposited with the CCDC and assigned CSD Number 2465609.